OpenStudy (anonymous):

use the distributive property to write the expression without parentheses 9(3+a)=

OpenStudy (anonymous):

would it be 9*3+a

OpenStudy (mathstudent55):

The distributive property is a(b + c) = ab + ac Multiply the number outside the parentheses by both numbers inside the parenthese and separate them by the operation inside the parentheses. Try again.

OpenStudy (anonymous):

27a

OpenStudy (anonymous):

or 27+a

OpenStudy (mathstudent55):

a(b + c) = ab + ac 9(3 + a) Multiply the 9 by 3: 27 Write the operation inside the parentheses: 27 + Multiply the 9 by 3: 27 + 9a

OpenStudy (anonymous):

36

OpenStudy (mathstudent55):

It's called the _distributive_ property because you're "distributing" what is outside the parentheses by everything inside the parenthsess

OpenStudy (mathstudent55):

27 and 9a cannot be added together because they are not like terms. The answer is 27 + 9a